OpenStudy (anonymous):

What is the numerator of the fraction equivalent to 4/15 that has a denominator of 60? (Points : 1) 6 8 12 16

OpenStudy (kropot72):

The question can be put as an equation: \[\large \frac{4}{15}=\frac{x}{60}\] The equation can be solved easily from the fact that the denominator on the right hand side is four times the value of the denominator on the left hand side. Therefore x = 4 * 4.

jhonyy9 (jhonyy9):

do you understand it now ?

jhonyy9 (jhonyy9):

@NashGriersGirly